OpenStudy (anonymous):

The area of a circle is 28.26 square centimeters. What is its diameter (use 3.14 for ð)?

OpenStudy (ajprincess):

\[Area~of~a~circle=\frac{\pi d^2}{4}\] \[28.26=\frac{3.14*d^2}{4}\] Solve for d. Does that help? @preetyboiyntg